OpenStudy (anonymous):

Find the value of x so that the line passing through (6.-25) and (x,20) has a slope of -3. F. -9 G.0 H.9 J.11

OpenStudy (anonymous):

setup an equation: \[-3=\frac{ 20-(25) }{ x-6 }\] Now rearrange and solve for x \[x-6=\frac{ 20+25 }{ -3 }\] \[x=\frac{ 20+25 }{ -3 }+6\]

OpenStudy (anonymous):

Is the answer 9?

OpenStudy (anonymous):

check your signs

OpenStudy (anonymous):

-9?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no problem